OpenStudy (anonymous):

Integral dx/sqrt(1-2x-x^2) by trigonometric substitution

OpenStudy (kainui):

First off, I think it'd be a good idea to complete the square.

OpenStudy (anonymous):

2-(x+1)^2 am i right?

OpenStudy (kainui):

You got it, now make the substitution\[ (x+1)=\sqrt{2}\sin(\theta)\] because that will allow you to get rid of the square root on the denominator. I can explain further if you like, it's sort of weird, but just go through it like a u-substitution now where you find dx.

OpenStudy (anonymous):

it's not like... integral dx/sqrt 2-(x+1)^2... so x=sen-1 and dx=cos ???

OpenStudy (kainui):

See, by making this particular substitution, you'll have: \[\sqrt{2-2\sin^2(\theta)}\] which if you use this idenitity\[2\sin^2 \theta +2 \cos ^2 \theta =2\] it means that you'll have \[\sqrt{2\cos^2\theta}=\sqrt2 \cos \theta\] in the denominator.

OpenStudy (anonymous):

\[dx=\sqrt{2} \cos \theta \] ???

OpenStudy (kainui):

Yes, almost, just forgot the last part: \[dx=\sqrt2 \cos \theta d \theta\] Now just plug them in and you get a pretty nice integral. =)

OpenStudy (anonymous):

ohh thanks i really appreciate :)